OpenStudy (anonymous):

@unknownunknown Sarah opened a simple interest savings account with $4000 in principal and an interest rate of 3.3%. How much is in the account after 6 years? simple interest: I = P • r • t

OpenStudy (anonymous):

@unknownunknown

OpenStudy (unknownunknown):

P = 4000, r= 0.033, t = 6 P*r*t = 4000*0.033 *6 + 4000 = 4792